# Wormhole Formation with Gravitational Particle Creation Mechanism

### v1

@@ II Wormhole Solution from Particle Creation Mechanism
In a homogeneous and isotropic Universe, the volume of the Universe may be represented using the scale factor $a$ as $V=a^{3}(t)$. Equation (2) may be recast in terms of total time derivatives of physical quantities as follows [25]:
@@
. As homogenous systems do not receive heat, assuming the Universe is homogeneous, we can speculate that the heat received remains constant over time ($dQ/dt=0$). This result enables us to reformulate equation (4) in an equivalent form:
@@
where $\rho(r)$ is the energy density, $p_{r}(r)$ and $p_{t}(r)$ are the radial and tangential pressures respectively which are mutually orthogonal. $\Pi(r)$ is the particle creation pressure introduced into the energy-momentum tensor $T^{\mu}_{\nu}=\text{diag}[-\rho(r),p_{r}(r)+\Pi(r),p_{t}(r)+\Pi(r),p_{t}(r)+\Pi(r)]$. Subsituting $\rho$ and (9) into equation (8) we obtain
@@
Now, substituting equation (12) in equation (11), we obtain the shape function for our model as
@@
The shape function given by equation (13) determines the configuration of the wormhole which is sustained by the negative particle creation pressure. Interestingly, the wormhole does not need an ad-hoc introduction of exotic fluids to account for the negative pressure required to sustain the wormhole geometry. To explore the physical nature of the matter content inside the wormhole geometry, we shall make use of the energy conditions. A generic aspect of wormhole systems is the violation of the null energy condition (NEC) (i.e. $\rho+p_{r}\geq 0$ or $\rho+p_{t}(r)\geq 0$ in terms of radial and tangential pressures respectively) at the throat of the wormhole structure. On the other hand, for some specific models keeping valid the NEC while keeping the dominant energy condition (DEC) ($\rho-p_{r}(r)\geq 0$ or $\rho-p_{t}(r)\geq 0$) violated at the throat is also possible [35, 36, 37, 38, 39]. In this work, we shall construct traversable wormhole structures with the violation of the NEC as well as DEC at the throat of the wormhole, simultaneously keeping the energy condition $\rho\geq 0$. This turns out, as a result of the violation of the weak energy condition (WEC) ($\rho\geq 0,\rho+p_{r}(r)\geq 0\text{ or }\rho+p_{t}(r)\geq 0$). From equation (10), given the shape function (13), one may obtain the energy density, radial and tangential pressures given by
@@ III.1 Shape function 1: $b(r)=\frac{r_{0}\tanh(\epsilon r)}{\tanh(\epsilon r_{0})}$
The particle creation pressure for the shape function (21) can be obtained from equation (11)
@@
From Fig. 4, it is observed that the NEC is violated in terms of radial pressure shown at the throat but tends to obey at the asymptotic limit, whereas the DEC is obeyed at the throat as well as at the asymptotic limit. In terms of the tangential pressure, the NEC is obeyed at the throat as well as at asymptotic regions but DEC is obeyed at the throat for the parameter values $(\beta,\epsilon)=(0.2,0.10)\text{ and }(0.5,0.15)$, however, crosses into the violation region at around $r=3$ and $r=1.6$ respectively for the ranges we considered. For $(\beta,\epsilon)=(0.7,0.20)$, the DEC is always violated. Also, as $\rho\geq 0$, WEC is violated for the radial pressure case and obeyed for the tangential pressure one. Moreover, from equation (25) and Fig. 5, it is evident that the particle creation pressure is negative at the throat all the way to $r\to\infty$, which is required and expected for backreaction pressure responsible for sustaining a wormhole geometry.
@@ III.2 Shape function 2: $b(r)=r_{0}^{n}r^{1-n}$
Fig. 7, shows that the NEC is violated in terms of radial pressure at the throat but tends to obey as $r\to\infty$, whereas the DEC is obeyed at the throat as well as at $r\to\infty$. In terms of the tangential pressure, the NEC is obeyed at the throat as well as at asymptotic regions but DEC is violated at the throat and obeyed at $r\to\infty$. Also, WEC is violated in terms of the radial pressure but obeyed in terms of the tangential pressure. Moreover, from equation (34) and Fig. 8 (left), it is seen that the particle creation pressure is negative at the throat to $r\to\infty$, but with an exception for the shape function (39) in that the particle creation pressure does not tend to 00 as $r\to\infty$ like all the other cases, but tends to higher negative values at regions far away from the throat.
@@ III.3 Shape function 3: $b(r)=r\exp[\Omega(r-r_{0})]$
From Fig. 10, we see NEC violation in terms of radial pressure at the throat but is obeyed at $r\to\infty$, whereas the DEC is obeyed at the throat as well as at $r\to\infty$. In terms of the tangential pressure, the NEC is satisfied at the throat as well as at asymptotic regions but DEC remains violated at the throat but obeyed at $r\to\infty$. Moreover, WEC is violated for the radial pressure case and satisfied for the tangential pressure one. Also, from equation (43) and Fig. 8 (right), it is seen that the particle creation pressure is negative at the throat and tends to zero at $r\to\infty$.
@@
which can now be compared with equation (48) to obtain
@@
Equation (51) can be solved to obtain a functional form of $z(r)$. The shape function obtained from the physical requirements of particle creation given by (13) is shown in the Fig. 11 for $\beta=0.7$ and $K=0.016$.

### v2

@@ II Wormhole Solution from Particle Creation Mechanism
In a homogeneous and isotropic Universe, the volume of the Universe may be represented using the scale factor $a$ as $V=a^{3}(t)$. Eq. (2) may be recast in terms of total time derivatives of physical quantities as follows [25]:
@@
. As homogenous systems do not receive heat, assuming the Universe is homogeneous, we can speculate that the heat received remains constant over time ($dQ/dt=0$). This result enables us to reformulate Eq. (4) in an equivalent form:
@@
where $\rho(r)$ is the energy density, $p_{r}(r)$ and $p_{t}(r)$ are the radial and tangential pressures respectively which are mutually orthogonal. $\Pi(r)$ is the particle creation pressure introduced into the energy-momentum tensor $T^{\mu}_{\nu}=\text{diag}[-\rho(r),p_{r}(r)+\Pi(r),p_{t}(r)+\Pi(r),p_{t}(r)+\Pi(r)]$. Subsituting $\rho$ and (9) into Eq. (8) we obtain
@@
Now, substituting Eq. (12) in Eq. (11), we obtain the shape function for our model as
@@
The shape function given by Eq. (13) determines the configuration of the wormhole which is sustained by the negative particle creation pressure. Interestingly, the wormhole does not need an ad-hoc introduction of exotic fluids to account for the negative pressure required to sustain the wormhole geometry. To explore the physical nature of the matter content inside the wormhole geometry, we shall make use of the energy conditions. A generic aspect of wormhole systems is the violation of the null energy condition (NEC) (i.e. $\rho+p_{r}\geq 0$ or $\rho+p_{t}(r)\geq 0$ in terms of radial and tangential pressures respectively) at the throat of the wormhole structure. On the other hand, for some specific models keeping valid the NEC while keeping the dominant energy condition (DEC) ($\rho-p_{r}(r)\geq 0$ or $\rho-p_{t}(r)\geq 0$) violated at the throat is also possible [35, 36, 37, 38, 39]. In this work, we shall construct traversable wormhole structures with the violation of the NEC as well as DEC at the throat of the wormhole, simultaneously keeping the energy condition $\rho\geq 0$. This turns out, as a result of the violation of the weak energy condition (WEC) ($\rho\geq 0,\rho+p_{r}(r)\geq 0\text{ or }\rho+p_{t}(r)\geq 0$). From Eq. (10), given the shape function (13), one may obtain the energy density, radial and tangential pressures given by
@@ III.1 Shape function 1: $b(r)=\frac{r_{0}\tanh(\epsilon r)}{\tanh(\epsilon r_{0})}$
The particle creation pressure for the shape function (21) can be obtained from Eq. (11)
@@
From Fig. 4, it is observed that the NEC is violated in terms of radial pressure shown at the throat but tends to obey at the asymptotic limit, whereas the DEC is obeyed at the throat as well as at the asymptotic limit. In terms of the tangential pressure, the NEC is obeyed at the throat as well as at asymptotic regions but DEC is obeyed at the throat for the parameter values $(\beta,\epsilon)=(0.2,0.10)\text{ and }(0.5,0.15)$, however, crosses into the violation region at around $r=3$ and $r=1.6$ respectively for the ranges we considered. For $(\beta,\epsilon)=(0.7,0.20)$, the DEC is always violated. Also, as $\rho\geq 0$, WEC is violated for the radial pressure case and obeyed for the tangential pressure one. Moreover, from Eq. (25) and Fig. 5, it is evident that the particle creation pressure is negative at the throat all the way to $r\to\infty$, which is required and expected for backreaction pressure responsible for sustaining a wormhole geometry.
@@ III.2 Shape function 2: $b(r)=r_{0}^{n}r^{1-n}$
Fig. 7, shows that the NEC is violated in terms of radial pressure at the throat but tends to obey as $r\to\infty$, whereas the DEC is obeyed at the throat as well as at $r\to\infty$. In terms of the tangential pressure, the NEC is obeyed at the throat as well as at asymptotic regions but DEC is violated at the throat and obeyed at $r\to\infty$. Also, WEC is violated in terms of the radial pressure but obeyed in terms of the tangential pressure. Moreover, from Eq. (34) and Fig. 8 (left), it is seen that the particle creation pressure is negative at the throat to $r\to\infty$. Similar observation is drawn for the shape function (39), given by Eq. (43) and shown in Fig. 8 (right) the particle creation pressure tends to zero as $r\to\infty$ like all the other cases.
@@ III.3 Shape function 3: $b(r)=r\exp[\Omega(r-r_{0})]$
From Fig. 10, we see NEC violation in terms of radial pressure at the throat but is obeyed at $r\to\infty$, whereas the DEC is obeyed at the throat as well as at $r\to\infty$. In terms of the tangential pressure, the NEC is satisfied at the throat as well as at asymptotic regions but DEC remains violated at the throat but obeyed at $r\to\infty$. Moreover, WEC is violated for the radial pressure case and satisfied for the tangential pressure one. Also, from Eq. (43) and Fig. 8 (right), it is seen that the particle creation pressure is negative at the throat and tends to zero at $r\to\infty$.
@@
which can now be compared with Eq. (48) to obtain
@@
Eq. (51) can be solved to obtain a functional form of $z(r)$. The shape function obtained from the physical requirements of particle creation given by (13) is shown in the Fig. 11 for $\beta=0.7$ and $K=0.016$.